1- A mixture of hydrogen and argon gases, in a 8.69 L flask at 77 °C, contains 0.477 grams of hydrogen and 7.10 grams of argon. The partial pressure of argon in the flask is ____ atm and the total pressure in the flask is ____ atm.

2- A mixture of methane and carbon dioxide gases is maintained in a 7.86 L flask at a pressure of 1.93 atm and a temperature of 39 °C. If the gas mixture contains 5.65 grams of methane, the number of grams of carbon dioxide in the mixture is ____ g.

3- The stopcock connecting a 2.65 L bulb containing oxygen gas at a pressure of 8.75 atm, and a 5.41 L bulb containing carbon dioxide gas at a pressure of 2.80 atm, is opened and the gases are allowed to mix. Assuming that the temperature remains constant, the final pressure in the system is ____ atm.

4- A mixture of helium and carbon dioxide gases contains helium at a partial pressure of 205 mm Hg and carbon dioxide at a partial pressure of 760 mm Hg. What is the mole fraction of each gas in the mixture?

XHe =
XCO2 =

5- A mixture of nitrogen and methane gases at a total pressure of 694 mm Hg contains nitrogen at a partial pressure of 517 mm Hg. If the gas mixture contains 7.04 grams of nitrogen, how many grams of methane are present?

____ g CH4
Answer #1

hole of H = 0.47-0.9385 mall q mole of Ar= = 0.174 mole 7.10 39.948 total mole=0.4155 mole Prosal (total poussere] = nx RXT =Total moles mole of O, + mole of Co, 9:15*2.65 + 2.8075.41 - RXT RXT Princex (2:6545.41 RXT Prince = 4.456 atm Ptotal = pre +mea of Cr, - 0.086 069 gm of CHg = mole x molar mass = 0.086069 x16 = 1.377 gm